OpenStudy (anonymous):

Math help please ?? Rewrite with only sin x and cos x. cos 3x the answer choices are : a.cos x - 4 cos x sin2x b.-sin3x + 2 sin x cos x c. -sin2x + 2 sin x cos x d.2 sin2x cos x - 2 sin x cos x

OpenStudy (yanasidlinskiy):

Hi Tofu! Here is your answer: cos(3x) = cos(x+x+x) = cos(x+2x) = cos(x)*cos(2x) - sin(x)*sin(2x) = cos(x)*(cos(x)*cos(x) - sin(x)*sin(x)) - sin(x)*(sin(x)*cos(x)+cos(x)*sin(x)) =cos^3(x) - sin^2(x)*cos(x) - sin^2(x)*cos(x) - sin^2(x)*cos(x) = cos^3(x) - 3*sin^2(x)*cos(x)

OpenStudy (anonymous):

i got that but its not any of the answer choices that's why im confused

OpenStudy (yanasidlinskiy):

@phi

OpenStudy (phi):

trig expressions can be manipulated into many different forms. It might be easier to test a few angles and see which of the choices matches with cos(3x) at that angle.

zepdrix (zepdrix):

Can you list the choices maybe? :)

OpenStudy (anonymous):

@zepdrix

zepdrix (zepdrix):

And welcome to OpenStudy! ^^

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

a.cos x - 4 cos x sin2x b.-sin3x + 2 sin x cos x c. -sin2x + 2 sin x cos x d.2 sin2x cos x - 2 sin x cos x

zepdrix (zepdrix):

Oh you did list them :P My bad.

OpenStudy (anonymous):

those are the answer choices

OpenStudy (anonymous):

it's ok

OpenStudy (anonymous):

its ok

zepdrix (zepdrix):

So you were able to get it to this point?\[\Large\rm \cos^3x - 3\sin^2x\cos x \]

zepdrix (zepdrix):

Oh boy, the site is tweaking out :(

OpenStudy (anonymous):

oh, dear but yes i dont know what to do after that

zepdrix (zepdrix):

Hmm looks like all of your options are in terms of 1st powers. So let's turn to our Half-Angle Identities.

OpenStudy (anonymous):

ok

zepdrix (zepdrix):

Start by breaking up the cos^3 into cos and cos^2 \[\Large\rm \cos x\color{orangered}{\cos^2x} - 3\color{royalblue}{\sin^2x}\cos x\]

zepdrix (zepdrix):

Then we can plug in our half angles,\[\Large\rm \color{orangered}{\cos^2x=\frac{1}{2}(1+\cos2x)}\]Remember the one for sine? :) It's very similar?

OpenStudy (anonymous):

1/2 (1- cos 2x) right ?

zepdrix (zepdrix):

mhm :)

zepdrix (zepdrix):

\[\Large\rm \cos x\color{orangered}{\cdot\frac{1}{2}(1+\cos2x)} - 3\color{royalblue}{\cdot\frac{1}{2}(1-\cos2x)}\cos x\]So you'll have to expand some stuff out and combine like-terms it appears.

OpenStudy (anonymous):

do i distribute both the 3 and cos x?

zepdrix (zepdrix):

In the second term? Looks like you'll need to distribute the 3, the cosx, the 1/2 AND the negative sign! :)

OpenStudy (anonymous):

ughh ok

OpenStudy (anonymous):

ahhhhhh idk if i did this right >.<

zepdrix (zepdrix):

XD

OpenStudy (anonymous):

ok so far i got \[(\cos x \div2+\cos ^{2}2x \div2)\]

OpenStudy (anonymous):

forr the first part

zepdrix (zepdrix):

Uh oh I think we're going down the wrong path :( See how all of the options have either sin3x or sin2x in them? I think maybe we weren't supposed to expand aaaall the way down to x's.

OpenStudy (anonymous):

oh no >.< i asked my teacher but she was no help to me

zepdrix (zepdrix):

\[\Large\rm \cos 3x=\cos(2x+x)=\cos(2x)\cos x-\sin(2x)\sin x\]Hmm none of the options have sin(2x)sinx, so I guess we can't leave that term alone.....

zepdrix (zepdrix):

This is a such a stupid question to pose as multiple choice :C ugh...

zepdrix (zepdrix):

I dunno >:c I give up, I need a break for now lol

OpenStudy (anonymous):

THAT'S I WAS SAYING -.-!!

OpenStudy (anonymous):

lol but thanks for the help :)

zepdrix (zepdrix):

Hmm I got an answer very close to option A. Lemme show ya the steps.

zepdrix (zepdrix):

\[\Large\rm \cos 3x=\cos(2x+x)=\cos(2x)\cos x-\sin(2x)\sin x\]Expanding cos(2x) and sin(2x) eventually gets us to: \[\Large\rm =\cos^3x - 3\sin^2x\cos x\]Which is what you had before, Let's steal 2 cosines from the first term:\[\Large\rm =\cos x(\cos^2x) - 3\sin^2x\cos x\]We'll use our square identity to write it as 1-sin^2x,\[\Large\rm =\cos x(1-\sin^2x) - 3\sin^2x\cos x\]Expanding and combining like-terms gives us,\[\Large\rm =\cos x-4\cos x \sin^2x\]We'll seperate the 2 sines,\[\Large\rm =\cos x-4(\cos x \sin x)\sin x\]Oh I guess we also need to steal 2 from that 4,\[\Large\rm =\cos x-2(2\sin x\cos x)\sin x\]And apply our Double Angle Formula to the bracketed part,\[\Large\rm =\cos x-2(\sin2x)\sin x\]

zepdrix (zepdrix):

We end up with:\[\Large\rm \cos x-2 \sin x \sin2x\]And option A is:\[\Large\rm \cos x-4\sin x \sin 2x\] Soooooo.. yah that's frustrating.